OpenStudy (impracticaljoker):

[02.03] Solve for x: −5|x + 1| = 10 (1 point) x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

Nnesha (nnesha):

to deal with absolute value first get rid of any number which is outside absolute value \[\left| \right|\] so first divide both sides by-5 what would you get on left and right side ??

OpenStudy (impracticaljoker):

I thought that I would add 5 to both sides?

Nnesha (nnesha):

good point! \[-5\left| x+1 \right|=10\] is same as \[-5 * \left| x+1 \right|=10\] there is multiplication..therefore we should do opposite of multiply to get rid of -5

Nnesha (nnesha):

if it was \[-5 + \left| x+1 \right|=10\] then you would add 5 both sides

OpenStudy (impracticaljoker):

Oh, I for what the | lines mean

OpenStudy (impracticaljoker):

|x + 1|= 2?

Nnesha (nnesha):

yeah thats the absolute value sign for example \[\left| x \right| \] the absolute value of x

Nnesha (nnesha):

what is 10 divide by -5 = ??

OpenStudy (impracticaljoker):

-2

Nnesha (nnesha):

\[\frac{-5 \left| x+1 \right| }{ -5 }=\frac{10}{-5}\]

OpenStudy (impracticaljoker):

wait waaa

OpenStudy (impracticaljoker):

i thought that was just to get rid of the -5

Nnesha (nnesha):

correct the absolute value of a number would always be positive there is some rule to solve absolute value \[ \left| x \right| = k\] where `k` is constant which should be positive \[k \ge 0\]

OpenStudy (impracticaljoker):

Wait, Im confused.

Nnesha (nnesha):

`i thought that was just to get rid of the -5` right so -5/-5 is just one \[\frac{ \cancel{-5 }\left| x+1 \right| }{\cancel{ -5} } = \frac{10}{-5}\] \[\left| x+1 \right|=-2\] i was just trying to write each step ..

Nnesha (nnesha):

here is an example \[\left| x \right| =- k\] where k is constant then the equation has no solution

Nnesha (nnesha):

\[\left| a+5 \right| = -9 ~~~ ~~~~~~~\left| x^2+8 \right|=-5\] more e xamples the equations doesn't have any solution you can't solve from there because there is a negative side to the opposite side of the absolute value

Nnesha (nnesha):

\[ \rm \left| x \right| = k \] (k is constant) in order to find the solution that k must be positive otherwise no solution

OpenStudy (phi):

after you divide both sides by -5, you get | x+1|= -2 the left side | x+1 | means "whatever x+1 is, make it positive" for example, if x was -2: | -2 + 1 | becomes | -1| and the | | means make the -1 a +1 so |-2+1| becomes (is the same as) 1 however, notice this: no matter what x we put inside | x+1 | the | | will make the answer positive. in other words, there is no way to make |x+1| equal a -2 so there is no solution i.e. no x that makes this equation true.